OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (jiteshmeghwal9):

1.) y=x2+2x-x+2 y=x(x+2)-1(x+2) y=(x-1)(x+2)

OpenStudy (jiteshmeghwal9):

in first equation (x+2) is common :)

OpenStudy (anonymous):

Solution of the equations: 1) y = x2 + 3x +2 D = 9 - 8 = 1 x1 = (-3+1)/2=-1 x2 = (-3-1)/2 = -2 2)y = x2 +2x +1 D = 4 - 4 = 0 x1 = x2 = -2/2 = -1 The solution means that the first graph has two points on the X axis, the second only one.

OpenStudy (anonymous):

Quadratic equations have anything in common, don't understand that question. Well, they both have solutions, and they are integers. The first one has two, while the second only one. Coefficients are different. Kind weird, never had such questions where I live.

OpenStudy (anonymous):

aha, well that's Florida for ya.